OpenStudy (anonymous):

Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + 3y inches, the other side should be x^2+7xy-9y^2 inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution.

OpenStudy (anonymous):

(x+3y)(x^2+7xy-9y^2) =(x^3+7x^2y- 9xy^2+3x^2y+21xy^2-27y^3) =x^3-27y^3 +10x^2y+12xy^2 square inches